A precipitation reaction occurs when 50.0 ml of 0.50 M BaBl2 is mixed with 75.0 mL of 0.75 M Na2CO3. The only precipitate is BaCO3 formed.


Write Balanced Equation

Limiting reactant

Theoretical Mass oF BaCO3

Actual Mass of BaCO3


1
Expert's answer
2022-01-03T17:03:03-0500

Balanced and Complete Reaction.

Na2CO3 + BaCl2 = BaCO3 + 2 NaCl .


Milli Moles of BaCl2 = ( 50 × 0.5 ) = 25

Milli Moles of Na2CO3 = ( 75 × 0.75 ) = 56.25


So, BaCl2 have less amount So, BaCl2 is limiting reagent .


Mass of BaCl2 = 25 × 208.23 = 5205.75 mg .

Mass of BaCl2 = 5.20575 g .


If 100 % yield then 5.20575 g is actual mass


If yield is smaller than then actual mass is smaller than theoritical .